OpenStudy (andijo76):

budget(x) 63,87,49,31,193,96,91 gross(y) 61,68,45,62,536,142,48 budget (x) =63,87,49,31,193,96,91=610 gross(y)= 61,68,45,62,536,142,48=962 x*y= 3843,5916,2205,1922,103448,13632,4368=130966 x^2= 3962,7569,2401,961,37249,9216,8281=70608 y^2= 3721,4624,2025,3844,287296,20164,2304=323978 i will show what i have so far

OpenStudy (andijo76):

numerator of r=(130966)-(610)(962)/7= 47134.571428 denominator of r= SQRT [7068-(610)^2/7]*SQRT[323978(962)^2/7] =Sqrt[17450.857142] =Sqrt[191771.714285] r=17450.857142 now I am stuck

OpenStudy (anonymous):

what math is this!!!????

OpenStudy (agreene):

What are you trying to do exactly? You didn't really give any context as to what you're doing.

OpenStudy (andijo76):

its statistics

OpenStudy (andijo76):

i was just taking a chance to see if anyone in here could answer this

OpenStudy (agreene):

Well, you haven't asked a question. So, it's hard to help.

OpenStudy (andijo76):

ok i will post the question sorry lol sometimes i space things off

OpenStudy (andijo76):

Listed below are the budgets (in millions of dollars and the gross receipts (in millions of dollars) for randomly selected movies construct a scatterplot find the value of the linear correlation coefficient r and find the P value using a=0.05 is there sufficient evidence to conclude that there is a linear correlation between budgets and gross receipts do the results change if the actual budgets listed are $63,000.000, $87,000.000, 49,000.000 and so on the numbers again are Budget (x) 63 87 49 31 193 96 91 Gross (y) 61 68 45 62 536 142 48